# Point Neuron Learning: A New Physics-Informed Neural Network Architecture

### v1

@@ 2 Problem formulation
Consider a source-free region of space denoted by $\Omega\subseteq\mathbb{R}^{2\mathrm{or}3}$ surrounded by sound sources as shown in Fig. 1. The sound pressure at an arbitrary point in the Cartesian coordinate $\boldsymbol{x}=(x,y,z),\boldsymbol{x}\in\Omega$, can be represented by $P(\boldsymbol{x},k)\in\mathbb{C}$, where $k=2\pi f/c$ is the wave number, $f$ is the frequency, and $c$ is the speed of sound propagation. Let there are $Q$ number of observation points $\boldsymbol{x}_{q}\in\Omega$, $q=1,...,Q$, to measure the sound field, as shown in Fig. 1.
@@
$\underset{\boldsymbol{\mu}\in\mathbb{C}^{S}}{\arg\min}$
$\operatorname{s.t.}$ $\Delta^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})+k^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})=0$ (1a)
@@
[…]
@@
Then, the research problem becomes how to find the optimal model $\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})$ based on $Q$ observations that can minimize $\mathcal{L}$ as well as satisfy the wave propagation constrain in (1a).
@@ 3.1 Point neuron
The propagation of sound is governed by the Helmholtz wave equation (1a) which is a natural constraint of any valid sound field and limits the direct adaptation of existing neural network architectures to model the sound field. Therefore, to satisfy (1a), we embed the fundamental solution of the Helmholtz equation, i.e., the free space Green function, into a neural network architecture,
$G(\boldsymbol{x}|\boldsymbol{y},k)=\frac{e^{ik\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}}{4\pi\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}},$ (1b)
@@
$PN(\boldsymbol{x}|\boldsymbol{y},k)=\rVert\boldsymbol{y}\rVert_{2}e^{-ik\rVert\boldsymbol{y}\rVert_{2}}\frac{e^{ik\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}}{4\pi\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}.$ (1c)
@@
The $v^{\text{th}}$ point neuron unit has three inputs, $x$, $y$, and $z$ co-ordinates of the point $\boldsymbol{x}$, and three biases $B^{x}_{v},B^{y}_{v},\mathrm{and}~{}B^{z}_{v}\in\mathbb{R}$. In the top path, the three channels of inputs with biases first go through an activation function $(\cdot)^{2}$ individually and then are fully connected (with unity weights) to pass another activation function $\sqrt{(\cdot)}$ and multiplying with the wave number $k$ (can be considered as a constant weight) before passing through the final activation function $h^{(1)}_{0}(\cdot)$, which is the $0^{\text{th}}$ order spherical Hankel function of the first kind, given by
$h^{(1)}_{0}(r)=\frac{e^{ir}}{ir}.$ (1d)
In the bottom path, the three biases directly undergo the same process as in the top path, followed by the final activation function $1/h^{(1)}_{0}(\cdot)$ after multiplying the fixed weight $k$. Finally, the outputs of the top and bottom paths are multiplied together to generate the point neuron output. We construct the point neuron network by fully connecting $V$ point neuron units with neuron weights $w_{v}\in\mathbb{C}$, $v=1,\ldots,V$ as shown in Fig. 2. The point neuron and its weights are frequency-dependent.
@@ 3.2 Back propagation and training
$\underset{\boldsymbol{\mu}\in\mathbb{C}^{S}}{\arg\min}$ $\mathcal{L}=\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q},k;\boldsymbol{\mu})-P(\boldsymbol{x}_{q},k)|^{2}}+\lambda\rVert\boldsymbol{w}\rVert_{1},$ (1ea)
$\operatorname{s.t.}$ $\Delta^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})+k^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})=0$ (1eb)
@@
$w_{v}(n+1)=w_{v}(n)-\xi\frac{\partial\mathcal{L}(n)}{\partial w_{v}(n)^{*}},$ (1f)
@@
$\frac{\partial\mathcal{L}(n)}{\partial w_{v}^{*}(n)}=$ $\sum_{q=1}^{Q}\Big{(}\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\Big{)}\,\frac{D_{v}(n)}{D_{v}^{q}(n)}\,e^{-ik\big{(}D_{v}^{q}(n)-D_{v}(n)\big{)}}\,+\frac{1}{2}\lambda e^{i\theta_{v}(n)},$ (1g)
@@
$n$ is the iteration index of the training process, $\xi$ is the learning rate, $(\cdot)^{*}$ denotes complex conjugation, and $\theta_{v}(n)\in\mathbb{R}$ is the phases of complex weight $w_{v}$. The proof of (1g) is given in the Appendix A.
@@
$B_{v}^{\alpha}(n+1)=B_{v}^{\alpha}(n)-\xi\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{\alpha}(n)},$ (1h)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{x}(n)}=$ $\sum_{q=1}^{Q}~{}2\operatorname{\mathbb{R}e}\Bigg{\{}\big{(}\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big{)}^{*}~{}w_{v}~{}\frac{D_{v}(n)}{D_{v}^{q}(n)}~{}e^{ik\big{(}D_{v}^{q}(n)-D_{v}(n)\big{)}}~{}$ (1i)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{y}(n)}=$ $\sum_{q=1}^{Q}2\,\operatorname{\mathbb{R}e}\Bigg{\{}\big{(}\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big{)}^{*}~{}w_{v}~{}\frac{D_{v}(n)}{D_{v}^{q}(n)}~{}e^{ik\big{(}D_{v}^{q}(n)-D_{v}(n)\big{)}}~{}$ (1j)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{z}(n)}=$ $\sum_{q=1}^{Q}~{}2\operatorname{\mathbb{R}e}\Bigg{\{}\big{(}\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big{)}^{*}~{}w_{v}~{}\frac{D_{v}(n)}{D_{v}^{q}(n)}~{}e^{ik\big{(}D_{v}^{q}(n)-D_{v}(n)\big{)}}~{}$ (1k)
@@
where $\operatorname{\mathbb{R}e}\{\cdot\}$ represents the real part of the argument. The proofs of (1i), (1j) and (1k) are given in Appendix B.
@@
    From (1g), (1i), (1j), and (1k), gradients of $w_{v}(n)$, $B_{v}^{x}(n)$, $B_{v}^{y}(n)$, and $B_{v}^{z}(n)$ become infinity when $D_{v}(n)$ and $D_{v}^{q}(n)$ becomes zero, which means virtual sources cannot be located at observer point positions or the origin point. Considering the physical meaning of the network, observer points cannot be on the source positions. Thus, during the training process, we apply a strategy to avoid placing virtual sources on the observer positions nor at the origin. We calculate the $D_{v}(n)$ and $D_{v}^{q}(n)$ in each iteration and replace $B_{v}^{x}(n)$, $B_{v}^{y}(n)$, and $B_{v}^{z}(n)$ to another location when $D_{v}(n)$ or $D_{v}^{q}(n)$ is smaller than a certain threshold.
@@ 3.3 System initialization
We initialise point neuron weights $w_{v}$ such that $|w_{v}|$ in the range of $[-1,1]$. For the initialization of system biases, we use any Prior-known knowledge about the underlying physical scenario to determine the initial positions of the virtual sources. For example, for modelling the outgoing sound field of a certain sound source, such as a drone, the virtual sources can be initially placed within a sphere that encompasses the sound source. The initialization will be further discussed in the next section.
@@ 4.1 Room settings
We evaluate the proposed method in a rectangular reverberant room with dimensions $6.0$ m $\times$ $4.0$ m $\times$ $4.0$ m where the coordinate origin is in the middle of the room. The target reconstruction region $\Omega$ is a horizontal circular region with a radius of $R=1.0$ m, centred at the point $(-1.0,0.5,0.0)$ m. There are five sources with equal strength (set to unity) located at the same horizontal plane, with coordinates $(-2.65,1.5,0.0)$ m, $(-2.4,-1.2,0.0)$ m, $(0.2,-1.5,0.0)$ m, $(1.7,-0.2,0.0)$ m, and $(1.0,1.2,0.0)$ m. We use the image source method [54] with wall reflection coefficients $[0.8,0.8,0.8,0.8,0,0]$ to simulate the sound field over the target reconstruction region. The corresponding reverberation time $RT_{60}$ is $0.537$ s [55]. Note that, we set the reflection coefficients of the roof and ground to zero to simplify the illustration.
[…]
@@ 4.2 Network settings
We randomly initialize the network weights by $|w_{v}|\in[-1,1]$. For the system biases, we initially place them over a mesh grid on the same horizontal plane as $\Omega$, excluding the $\Omega$, with a size of 9.0 m $\times$ 9.0 m. The number of point neurons increases from $V=25$ for the lowest frequency to $V=465$ for the highest frequency. The parameter $\lambda$ is in the range of $[3\times 10^{-4},3\times 10^{-3}]$, and $\xi=3\times 10^{-2}$.
We compare the proposed sound field reconstruction with a conventional method and a typical PINN method. The conventional method is the orthogonal harmonics-based sound field estimation method that reconstructs any incident sound field by a finite number of harmonic basis functions and corresponding coefficients [57, 50]. We adopt the network architecture proposed in [44] as the typical PINN method since it adds PDE loss to a fully connected network architecture similar to many PINN methods. Similar to the proposed method, the PINN in [44] also only requires microphone observations in the training process. This network contains 3 layers and 15 nodes of each layer, with the activation function being $tanh$, and initializes the trainable parameters with the Xavier initialization. We train the PINN for $108$ epochs with a learning rate of $10^{-5}$ using the ADAM optimizer. The data loss of the network is calculated by all microphones, and the PDE loss is calculated by $437$ uniformly arranged sampling points over the target region.
@@ 4.3 Evaluation metrics
$\mathrm{NMSE}(k)=20\mathrm{log}_{10}\frac{\sum^{M}_{m=1}|\mathcal{P}(\boldsymbol{x}_{m},k)-P(\boldsymbol{x}_{m},k)|}{\sum^{M}_{m=1}|P(\boldsymbol{x}_{m},k)|},$ (1l)
where $M$ is the number of evaluation points, $m=1,...,M$ be the index of evaluation points, and $\boldsymbol{x}_{m}\in\Omega$.
@@
[…]
@@
$\mathrm{NSE}(\boldsymbol{x},k)=20\mathrm{log}_{10}\frac{|\mathcal{P}(\boldsymbol{x},k)-P(\boldsymbol{x},k)|}{|P(\boldsymbol{x},k)|}.$ (1n)

### v2

@@ 2 Problem formulation
Consider a source-free region of space denoted by $\Omega\subseteq\mathbb{R}^{2\mathrm{or}3}$ surrounded by sound sources as shown in Fig. 1. The sound pressure at an arbitrary point in the Cartesian coordinate $\boldsymbol{x}=(x,y,z),\boldsymbol{x}\in\Omega$, can be represented by $P(\boldsymbol{x},k)\in\mathbb{C}$, where $k=2\pi f/c$ is the wave number, $f$ is the frequency, and $c$ is the speed of sound propagation. Let there are $Q$ number of observation points $\boldsymbol{x}_{q}\in\Omega$, q=1,,Qq=1,...,Q, to measure the sound field, as shown in Fig. 1.
@@
$\underset{\boldsymbol{\mu}\in\mathbb{C}^{S}}{\arg\min}$ $\mathcal{L}=\mathcal{L}_{\text{TRN}}(\boldsymbol{p}_{\mathrm{mic}},\mathbcal{p}_{\mathrm{mic}})+\lambda\mathcal{C}(\boldsymbol{\mu}),$ (1a)
$\operatorname{s.t.}$ $\Delta^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})+k^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})=0$ (1b)
@@
where $\mathcal{L}$ is the cost function, $\mathcal{L}_{\text{TRN}}$ is the training loss that measures the supervised error between the sound field estimation $\mathbcal{p}_{\mathrm{mic}}=[\mathcal{P}(\boldsymbol{x}_{1},k),\ldots,\mathcal{P}(\boldsymbol{x}_{Q},k)]^{T}$ and the pressure measurements $\boldsymbol{p}_{\mathrm{mic}}=[P(\boldsymbol{x}_{1},k),\ldots,P(\boldsymbol{x}_{Q},k)]^{T}$, $\lambda$ is a hyper-parameter to control the weight of model complexity loss $\mathcal{C}(\boldsymbol{\mu})$, and
@@
Then, the research problem becomes how to find the optimal model $\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})$ based on $Q$ observations that can minimize $\mathcal{L}$ as well as satisfy the wave propagation constrain in (1b).
@@ 3.1 Point neuron
The propagation of sound is governed by the Helmholtz wave equation (1b) which is a natural constraint of any valid sound field and limits the direct adaptation of existing neural network architectures to model the sound field. Therefore, to satisfy (1b), we embed the fundamental solution of the Helmholtz equation, i.e., the free space Green function, into a neural network architecture,
$G(\boldsymbol{x}|\boldsymbol{y},k)=\frac{e^{ik\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}}{4\pi\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}},$ (2)
@@
$PN(\boldsymbol{x}|\boldsymbol{y},k)=\rVert\boldsymbol{y}\rVert_{2}e^{-ik\rVert\boldsymbol{y}\rVert_{2}}\frac{e^{ik\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}}{4\pi\rVert\boldsymbol{x}-\boldsymbol{y}\rVert_{2}}.$ (3)
@@
[…]
$h^{(1)}_{0}(r)=\frac{e^{ir}}{ir}.$ (4)
[…]
@@ 3.2 Back propagation and training
$\underset{\boldsymbol{\mu}\in\mathbb{C}^{S}}{\arg\min}$ $\mathcal{L}=\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q},k;\boldsymbol{\mu})-P(\boldsymbol{x}_{q},k)|^{2}}+\lambda\rVert\boldsymbol{w}\rVert_{1},$ (5a)
$\operatorname{s.t.}$ $\Delta^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})+k^{2}\mathcal{P}(\boldsymbol{x},k;\boldsymbol{\mu})=0$ (5b)
@@
$w_{v}(n+1)=w_{v}(n)-\xi\frac{\partial\mathcal{L}(n)}{\partial w_{v}(n)^{*}},$ (6)
@@
$\frac{\partial\mathcal{L}(n)}{\partial w_{v}^{*}(n)}=$ $\sum_{q=1}^{Q}\Big(\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\Big)\,\frac{D_{v}(n)}{D_{v}^{q}(n)}\,e^{-ik\big(D_{v}^{q}(n)-D_{v}(n)\big)}\,+\frac{1}{2}\lambda e^{i\theta_{v}(n)},$ (7)
@@
$n$ is the iteration index of the training process, $\xi$ is the learning rate, $(\cdot)^{*}$ denotes complex conjugation, and $\theta_{v}(n)\in\mathbb{R}$ is the phases of complex weight $w_{v}$. The proof of (7) is given in the Appendix A.
@@
$B_{v}^{\alpha}(n+1)=B_{v}^{\alpha}(n)-\xi\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{\alpha}(n)},$ (8)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{x}(n)}=$ $\sum_{q=1}^{Q}~2\operatorname{\mathbb{R}e}\Bigg\{\big(\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big)^{*}~w_{v}~\frac{D_{v}(n)}{D_{v}^{q}(n)}~e^{ik\big(D_{v}^{q}(n)-D_{v}(n)\big)}$~ (9)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{y}(n)}=$ $\sum_{q=1}^{Q}2\,\operatorname{\mathbb{R}e}\Bigg\{\big(\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big)^{*}~w_{v}~\frac{D_{v}(n)}{D_{v}^{q}(n)}~e^{ik\big(D_{v}^{q}(n)-D_{v}(n)\big)}$~ (10)
@@
$\frac{\partial\mathcal{L}(n)}{\partial B_{v}^{z}(n)}=$ $\sum_{q=1}^{Q}~2\operatorname{\mathbb{R}e}\Bigg\{\big(\mathcal{P}(\boldsymbol{x}_{q},n)-P(\boldsymbol{x}_{q})\big)^{*}~w_{v}~\frac{D_{v}(n)}{D_{v}^{q}(n)}~e^{ik\big(D_{v}^{q}(n)-D_{v}(n)\big)}$~ (11)
@@
where $\operatorname{\mathbb{R}e}\{\cdot\}$ represents the real part of the argument. The proofs of (9), (10) and (11) are given in Appendix B.
@@
    From (7), (9), (10), and (11), gradients of $w_{v}(n)$, $B_{v}^{x}(n)$, $B_{v}^{y}(n)$, and $B_{v}^{z}(n)$ become infinity when $D_{v}(n)$ and $D_{v}^{q}(n)$ becomes zero, which means virtual sources cannot be located at observer point positions or the origin point. Considering the physical meaning of the network, observer points cannot be on the source positions. Thus, during the training process, we apply a strategy to avoid placing virtual sources on the observer positions nor at the origin. We calculate the $D_{v}(n)$ and $D_{v}^{q}(n)$ in each iteration and replace $B_{v}^{x}(n)$, $B_{v}^{y}(n)$, and $B_{v}^{z}(n)$ to another location when $D_{v}(n)$ or $D_{v}^{q}(n)$ is smaller than a certain threshold.
@@ 3.3 System initialization
We initialise point neuron weights $w_{v}$ such that $|w_{v}|$ in the range of [1,1][-1,1]. For the initialization of system biases, we use any Prior-known knowledge about the underlying physical scenario to determine the initial positions of the virtual sources. For example, for modelling the outgoing sound field of a certain sound source, such as a drone, the virtual sources can be initially placed within a sphere that encompasses the sound source. The initialization will be further discussed in the next section.
@@ 4.1 Room settings
We evaluate the proposed method in a rectangular reverberant room with dimensions $6.0$ m $\times$ $4.0$ m $\times$ $4.0$ m where the coordinate origin is in the middle of the room. The target reconstruction region $\Omega$ is a horizontal circular region with a radius of $R=1.0$ m, centred at the point (1.0,0.5,0.0)(-1.0,0.5,0.0) m. There are five sources with equal strength (set to unity) located at the same horizontal plane, with coordinates (2.65,1.5,0.0)(-2.65,1.5,0.0) m, (2.4,1.2,0.0)(-2.4,-1.2,0.0) m, (0.2,1.5,0.0)(0.2,-1.5,0.0) m, (1.7,0.2,0.0)(1.7,-0.2,0.0) m, and $(1.0,1.2,0.0)$ m. We use the image source method [54] with wall reflection coefficients $[0.8,0.8,0.8,0.8,0,0]$ to simulate the sound field over the target reconstruction region. The corresponding reverberation time $RT_{60}$ is $0.537$ s [55]. Note that, we set the reflection coefficients of the roof and ground to zero to simplify the illustration.
We consider two different microphone placements to capture the sound field: (i) uniform circular array on the boundary of the target region (Fig. 3a), and (ii) randomly selected locations inside the target region (Fig. 3b). We initially place $Q=75$ microphones based on $Q=2N+1$ where $N=\lceil{k_{f_{\text{max}}}R}\rceil$ and $\lceil{\cdot}\rceil$ is the ceiling operation [56]. In Sec. 4.6, we investigate the performance by varying the number of microphones. For all simulations, except Sec. 4.7, White Gaussian noise with an SNR of 20 dB is added to each of the microphone measurements.
@@ 4.2 Network settings
We randomly initialize the network weights by $|w_{v}|\in[-1,1]$. For the system biases, we initially place them over a mesh grid on the same horizontal plane as $\Omega$, excluding the $\Omega$, with a size of 9.0 m $\times$ 9.0 m. The number of point neurons increases from $V=25$ for the lowest frequency to $V=465$ for the highest frequency. The parameter $\lambda$ is in the range of $3\times 10^{-4},3\times 10^{-3}]$, and $\xi=3\times 10^{-2}$.
We compare the proposed sound field reconstruction with a conventional method and a typical PINN method. The conventional method is the orthogonal harmonics-based sound field estimation method that reconstructs any incident sound field by a finite number of harmonic basis functions and corresponding coefficients [57, 50]. We adopt the network architecture proposed in [44] as the typical PINN method since it adds PDE loss to a fully connected network architecture similar to many PINN methods. Similar to the proposed method, the PINN in [44] also only requires microphone observations in the training process. This network contains 3 layers and 15 nodes of each layer, with the activation function being tanhtanh, and initializes the trainable parameters with the Xavier initialization. We train the PINN for $108$ epochs with a learning rate of $10^{-5}$ using the ADAM optimizer. The data loss of the network is calculated by all microphones, and the PDE loss is calculated by $437$ uniformly arranged sampling points over the target region.
@@ 4.3 Evaluation metrics
$\mathrm{NMSE}(k)=20\mathrm{log}_{10}\frac{\sum^{M}_{m=1}|\mathcal{P}(\boldsymbol{x}_{m},k)-P(\boldsymbol{x}_{m},k)|}{\sum^{M}_{m=1}|P(\boldsymbol{x}_{m},k)|},$ (12)
where $M$ is the number of evaluation points, m=1,,Mm=1,...,M be the index of evaluation points, and $\boldsymbol{x}_{m}\in\Omega$.
@@
$\mathrm{MAC}(k)=\frac{\rVert\boldsymbol{p}_{\mathrm{eval}}^{H}\mathbcal{p}_{\mathrm{eval}}\rVert_{2}^{2}}{(\boldsymbol{p}_{\mathrm{eval}}^{H}\boldsymbol{p}_{\mathrm{eval}})(\mathbcal{p}_{\mathrm{eval}}^{H}\mathbcal{p}_{\mathrm{eval}})},$ (13)
where $\boldsymbol{p}_{\mathrm{eval}}=[P(\boldsymbol{x}_{1},k),\ldots,P(\boldsymbol{x}_{M},k)]^{T}$ and $\mathbcal{p}_{\mathrm{eval}}=[\mathcal{P}(\boldsymbol{x}_{1},k),\ldots,\mathcal{P}(\boldsymbol{x}_{M},k)]^{T}$ are two $M\times 1$ vectors that contain the original sound pressure and estimated sound pressure of evaluate points, $(\cdot)^{H}$ denotes the conjugate transpose.
@@
$\mathrm{NSE}(\boldsymbol{x},k)=20\mathrm{log}_{10}\frac{|\mathcal{P}(\boldsymbol{x},k)-P(\boldsymbol{x},k)|}{|P(\boldsymbol{x},k)|}.$ (14)